As an intern at Hadlock & Fee Investments, you have been assigned to manage a $10 million portfolio that is invested in five stocks. What is the portfolio's Beta?

                                    Stock               Amount Invested                    Beta

                                    TJX                 $4 million                                1.2

                                    GM                  $2 million                                1.1

                                    RTC                $2 million                                1.0

                                    MAB               $1 million                                0.7

INX                 $1 million                                0.5
Answer #1

А - B D E F G cu AwN- Stock TUX GM RTC MAB INX Total Investment Weight Beta 4000000 0.4 L 1.2 2000000 0.2 1.1 2000000 0.2 L 1

EXCEL FORMULA:

А B DE Beta Stock TJX GM RTC MAB INX Total Investment 4000000 2000000 2000000 1000000 1000000 =SUM(C3:07) Weight =C3/C8 1.2 =

FEEL FREE TO ASK DOUBTS... PLEASE RATE MY ANSWER